# Multi-Input Multi-Output Target-Speaker Voice Activity Detection For Unified, Flexible, and Robust Audio-Visual Speaker Diarization

### v2

@@ Abstract
Audio-visual learning has demonstrated promising results in many classical speech tasks (e.g., speech separation, automatic speech recognition, wake-word spotting). We believe that introducing visual modality will also benefit speaker diarization. To date, Target-Speaker Voice Activity Detection (TS-VAD) plays an important role in highly accurate speaker diarization. However, previous TS-VAD models take audio features and utilize the speaker’s acoustic footprint to distinguish his or her personal speech activities, which is easily affected by overlapped speech in multi-speaker scenarios. Although visual information naturally tolerates overlapped speech, it suffers from spatial occlusion, low resolution, etc. The potential modality-missing problem blocks TS-VAD towards an audio-visual approach.
This paper proposes a novel Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD) framework for speaker diarization. The proposed method can take audio-visual input and leverage the speaker’s acoustic footprint or lip track to flexibly conduct audio-based, video-based, and audio-visual speaker diarization in a unified sequence-to-sequence framework. Experimental results show that the MIMO-TSVAD framework demonstrates state-of-the-art performance on the VoxConverse, DIHARD-III, and MISP 2022 datasets under corresponding evaluation metrics, obtaining the Diarization Error Rates (DERs) of 4.18%, 10.10%, and 8.15%, respectively. In addition, it can perform robustly in heavy lip-missing scenarios.
@@ I Introduction
Like documenting events in a diary, speaker diarization is the task of automatically detecting multiple speakers’ utterance boundaries in conversational data [1]. It aims to split the audio or multi-modal signals into segments with labeled identities, solving the problem of ”Who-Spoke-When.” As a front-end technique, it is essential in various downstream applications (e.g., speech recognition) [2].
In previous studies, speaker diarization research mainly focuses on audio streams [2]. The modularized method, also known as the conventional method, utilizes cascaded modules to partition the audio signal into short segments and cluster their identities by advanced speaker representation techniques [3, 4, 5]. While, these methods cannot handle overlapped speech as each audio segment is supposed to be speaker-homogeneous. Some studies propose additional post-processing techniques (e.g., overlapped speech detection [6], overlap-aware resegmentation [7, 5]) to compromise this effect. Then, End-to-End Neural Diarization (EEND) systems [8, 9, 10, 11] are proposed to estimate multiple speakers’ speech activities as multi-label classification. The end-to-end structure of neural networks leads to ease of optimization and robustness to overlapped speech. Nevertheless, permutation-invariant training in EEND-based methods causes performance degradation when the number of speakers increases in long audios. Although a few studies [12, 13, 14, 15, 16] have explored the unsupervised clustering to address this problem, their results are still unsatisfactory. Recently, Target-Speaker Voice Activity Detection (TS-VAD) approaches [17, 18, 19, 20] become attractive, which combines advantage of modularized methods and end-to-end neural networks. As a post-processing method, the TS-VAD framework requires an initial diarization system (e.g., modularized method) to extract each speaker’s acoustic footprint as the speaker profile, solving the problem of ”Who-Spoke” in advance. Then, a neural network-based model takes speech features and all speaker profiles to predict their corresponding framewise voice activities, aiming to address the ”When” problem. This two-stage process has demonstrated excellent performance in popular benchmarks such as DIHARD-III [21] and VoxSRC21-23 [22, 23, 24].
@@
In this article, we propose a novel Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD) framework for audio-visual speaker diarization in modality-missing scenarios. Let $\mathbf{X}_{a}$ and $\mathbf{X}_{v}$ denote the audio and video features, respectively. $\mathbf{E}_{spk}$ represents the target-speaker embeddings in classical TS-VAD systems. We additionally define target-lip embeddings to indicate speaker identities of corresponding lip tracks, denoted as $\mathbf{E}_{lip}$. Depending on the accessibility of different data during inference, there are four cases , namely $\mathbf{X}_{a}$ vs. $\mathbf{E}_{spk}$, $\mathbf{X}_{v}$ vs. $\mathbf{E}_{lip}$, $\mathbf{X}_{a}+\mathbf{X}_{v}$ vs. $\mathbf{E}_{lip}$, $\mathbf{X}_{a}+\mathbf{X}_{v}$ vs. $\mathbf{E}_{spk}+\mathbf{E}_{lip}$. Our proposed framework can flexibly process arbitrary cases with state-of-the-art performance and strong robustness.
Fig. 1 illustrates the progress of our TS-VAD research. The classical TS-VAD systems [17, 18] can be abstracted into the encoder-only methods, shown in Fig. 1a. $\mathbf{X}_{a}\in\mathbb{R}^{\mathrm{T\times F}}$ is extracted from input audio by traditional methods (e.g., MFCC, Fbank) or deep neural networks. $\mathbf{E}_{spk}\in\mathbb{R}^{\mathrm{N\times S}}$ represents the given speaker embeddings. Each speaker embedding $\in\mathbb{R}^{\mathrm{S}}$ has to be repeated $T$ times and concatenated with $\mathbf{X}_{a}$, producing a 3-dimensional tensor with the shape of $T\times N\times(F+S)$. Then, encoders (e.g., Bi-LSTM, Transformers) process the input tensor along the time axis ($T$) and speaker axis ($N$). As shown in Fig. 1b, we introduce the sequence-to-sequence architecture [19]) to factorize the input onto the encoder and decoder separately. This way, the consumption of input memory becomes proportional to $T\times F+N\times S$, enabling the model to process longer audio and more speakers. Moreover, adjusting the output dimension ($T^{\prime}$) of the last linear layer can achieve a higher temporal resolution for predicted voice activities, regardless of the input feature length ($T$). In this work, the MIMO-TSVAD framework is designed based on the sequence-to-sequence architecture, demonstrated in Fig. 1c. The multi-modal encoder and multi-task decoder leverage cross-modal and inter-speaker relationships, which can utilize various types of input data to obtain different predictions in a Multi-Input and Multi-Output (MIMO) manner.
@@ II-B End-to-End Neural Diarization
End-to-End Neural Diarization (EEND) [8, 9] formulates the diarization problem as a multi-label classification task, relying on the permutation-invariant training to predict all speakers’ voice activities simultaneously. The initial EEND models have a fixed number of output speakers limited by the network architecture. Although the Encoder-Decoder based Attractor (EDA) [10, 11] enables EEND-based methods to process audio with a variable number of speakers, the maximum number of speakers is empirically bounded by the training data. To make the number of output speakers flexible and unlimited, EEND-vector clustering (EEND-VC) [12, 13, 14] integrates end-to-end and clustering approaches, which deploys an EEND model for shortly divided audio blocks and then matches the inter-block speaker labels by clustering on speaker embeddings. In addition, EEND-GLA [15, 16] calculates local attractors from each short block and finds the speaker correspondence based on similarities between inter-block attractors. As its training only requires relative speaker labels within the recording, EEND-GLA is practical for adapting models on in-the-wild datasets without globally unique speaker labels.
@@ II-C Target-Speaker Voice Activity Detection
The background of Target-Speaker Voice Activity Detection (TS-VAD) can be traced back to the personal VAD [61], which utilizes a given acoustic footprint as the speaker profile to retrieve his or her personalized voice activities. However, the personal VAD only accepts a single speaker at once and ignores the inter-speaker modeling in a conversation. Thus, TS-VAD is designed to process multiple speaker profiles and predict their voice activities simultaneously.
The initial TS-VAD [17] takes speech features (e.g., MFCC) and i-vectors of each speaker as the input, where the output number of speakers is fixed. He et al. [62] modify the model to cope with a flexible number of speakers by determining the maximum number of speakers and outputting null speech activities for zero-padded speaker profiles. Later, LSTM [63] and Transformer [20] modules are implemented along the speaker dimension of models to handle a variable number of speakers. On the other hand, the i-vectors used in TS-VAD are relatively domain-dependent, restricting the system performance on multi-scenario dataset [64]. This finding paves the way for exploring more discriminative speaker embeddings like x-vectors as an alternative. Wang et al. [18] first replace the front-end of TS-VAD with a pre-trained module tailored for extracting frame-level x-vectors. This modification shows superior robustness and generalization than a simple swap of i-vectors for x-vectors in early attempt [17].
@@
Figure 2: The Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD) framework. (a) Extractor: Front-end modules extract audio features $\mathbf{X}_{a}$ and video features $\mathbf{X}_{v}$ from the input data. (b): Encoder: Conformer-based modules leverage multi-modal information to generate the updated audio-visual features $\mathbf{\hat{X}}_{a}$ and $\mathbf{\hat{X}}_{v}$. (c) Decoder: Multi-task decoding modules predict voice activities based on different speaker profiles, including target-speaker embeddings $\mathbf{E}_{spk}$, target-lip embeddings $\mathbf{E}_{lip}$, or both. Green, blue, and yellow depict sub-components for processing audio, visual, and audio-visual modalities. For clarity, layer normalization and residual connection of each self-attention, cross-attention, and feedforward layer are omitted in the plot.
@@ II-D Audio-Visual Speaker Diarization
As facial attributes and lip movements have been proven to be highly related to speech [70], most early audio-visual speaker diarization methods leverage multi-modal cues by modeling the correspondence between speech signals and talking faces [25, 26]. Also, sound source localization using microphone arrays can establish another cross-modal relationship by mapping the speech direction onto the captured image plane [27, 28]. However, it is sometimes hard to perfectly enroll talking faces or lip movements with related speech segments. Off-screen speakers whose faces are not captured can lead to face detection failure. This ”enroll first, diarize later” paradigm may fail in modality-missing scenarios.
Recently, Xu et al. [29] create the AVA Audio-Visual Diarization (AVA-AVD) database containing diverse movie clips and a multi-stage audio-visual speaker diarization system. The proposed method first utilizes Voice Activity Detection (VAD) to find speaking segments in the audio, then applies the Active Speaker Detection (ASD) [34] to locate a face for each speaking segment. The paired audio-visual inputs are jointly scored to cluster the speaking segments into different identities. Subsequently, Wuerkaixi et al. [30] introduce the lip movement to learn a Dynamic Vision-guided Speaker Embedding (DyViSE) instead of the still facial feature, achieving new state-of-the-art performance on these real-world videos.
In addition, the Multi-Modal Information based Speech Processing (MISP) [71, 37] database is presented for Chinese Home-TV scenarios. It has over 100 hours of audio-visual signals synchronously captured by different devices, nearly 3.5 times larger than the AVA-AVD [29]. The MISP 2022 Challenge has been successfully held based on this database. Meanwhile, an End-to-End Audio-Visual Speaker Diarization (AVSD) [31] is presented as a competitive baseline method that predicts speech probabilities using audio features and each speaker’s lip video. In the AVSD system, available lip videos decide the number and order of output speakers. This way is highly effective when the camera captures all speakers ideally. In contrast, it cannot deal with the out-of-screen speakers.
@@ III-A Architecture
Fig. 2 demonstrates the framework our proposed Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD), consisting of three parts: the extractor, the encoder, and the decoder.
@@ III-A2 Video Extractor
The ResNet18-3D [73] is adopted as the video front-end extractor. A few modifications are employed based on its standard implementation in PytorchVideo [^1] https://github.com/facebookresearch/pytorchvideo. First, the stem layer is set to the convolutional kernel size of 7, stride of 2, and output channels of 32 without the pooling layer. The stride of pooling and convolutional layers in the residual blocks is set to $\left(1,2,2\right)$ without the temporal downsampling. Finally, a spatial global average pooling layer is placed at the tail. The model transforms a lip video with the length of $T_{2}$ and resolution of $H\times W$ into frame-level representations $\mathbf{X}_{v}\in\mathbb{R}^{\mathrm{T_{2}\times F_{2}}}$, where $F_{2}$ represents the feature dimension.
@@ III-A4 Decoder
The input decoder embeddings are initialized by zeros and updated step by step to learn temporal (positional) information about voice activities. Thus, the self-attention layer is adopted to exchange inter-speaker relationships. Then, $\mathbf{E}_{spk}$ and $\mathbf{E}_{lip}$ go through multi-layer perception (MLP) modules to align the feature dimension as same as the decoder size and concatenate with the original queries in respective branch. The adopted MLP module consists of two linear layers with in-between layer normalization and ReLU activation. Meanwhile, we concatenate keys from $\mathbf{\hat{X}}_{a}$ and $\mathbf{\hat{X}}_{v}$ with corresponding positional embeddings $\mathbf{P_{a}}$ and $\mathbf{P_{v}}$. This way, all key-query calculations in the cross-attention layer can incorporate positional and speaker-related information as the auxiliary feature. Finally, a linear projection layer with sigmoid activation transforms the decoder embeddings into posterior probabilities of the estimated voice activities. The output dimension $T^{\prime}$ of the linear projection controls the temporal resolution of system prediction. It can be set to several multiples of the feature length $T_{1}$ or $T_{2}$, which provides more precise timestamps. Given $N$ speakers, the predictions from the audio-based, video-based, and mixed output are represented as $\mathbf{\hat{Y}}_{aud}$, $\mathbf{\hat{Y}}_{vid}$, and $\mathbf{\hat{Y}}_{mix}\in\mathbb{R}^{\mathrm{N\times T^{\prime}}}$, respectively.

### v3

@@ Abstract
Audio-visual learning has demonstrated promising results in many classical speech tasks (e.g., speech separation, automatic speech recognition, wake-word spotting). We believe that introducing visual modality will also benefit speaker diarization. To date, Target-Speaker Voice Activity Detection (TS-VAD) plays an important role in highly accurate speaker diarization. However, previous TS-VAD models take audio features and utilize the speaker’s acoustic footprint to distinguish his or her personal speech activities, which is easily affected by overlapped speech in multi-speaker scenarios. Although visual information naturally tolerates overlapped speech, it suffers from spatial occlusion, low resolution, etc. The potential modality-missing problem blocks TS-VAD towards an audio-visual approach. This paper proposes a novel Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD) framework for speaker diarization. The proposed method can take audio-visual input and leverage the speaker’s acoustic footprint or lip track to flexibly conduct audio-based, video-based, and audio-visual speaker diarization in a unified sequence-to-sequence framework. Experimental results show that the MIMO-TSVAD framework demonstrates state-of-the-art performance on the VoxConverse, DIHARD-III, and MISP 2022 datasets under corresponding evaluation metrics, obtaining the Diarization Error Rates (DERs) of 4.18%, 10.10%, and 8.15%, respectively. In addition, it can perform robustly in heavy lip-missing scenarios.
@@ I Introduction
Like documenting events in a diary, speaker diarization is the task of automatically detecting multiple speakers’ utterance boundaries in conversational data [1]. It aims to split the audio or multi-modal signals into segments with labeled identities, solving the problem of “Who-Spoke-When.” As a front-end technique, it is essential in various downstream applications (e.g., speech recognition) [2].
In previous studies, speaker diarization research mainly focuses on audio streams [2]. The conventional method, also known as the modularized method, utilizes cascaded modules to partition the audio signal into short segments and cluster their identities by advanced speaker representation techniques [3, 4, 5]. These methods cannot handle overlapped speech as each audio segment is supposed to be speaker-homogeneous. Some studies propose additional post-processing techniques (e.g., overlapped speech detection [6], overlap-aware resegmentation [7, 5]) to compromise this effect. Then, End-to-End Neural Diarization (EEND) systems [8, 9, 10, 11] are proposed to estimate multiple speakers’ speech activities as multi-label classification. The end-to-end structure of neural networks leads to ease of optimization and robustness to overlapped speech. Nevertheless, permutation-invariant training in EEND-based methods causes performance degradation when the number of speakers increases in long audios. Although a few studies [12, 13, 14, 15, 16] have explored the unsupervised clustering to address this problem, their results are still unsatisfactory. Recently, TS-VAD approaches [17, 18, 19, 20] become attractive, which combine advantages of modularized methods and end-to-end neural networks. As a post-processing method, the TS-VAD framework requires an initial diarization system (e.g., modularized method) to extract each speaker’s acoustic footprint as the speaker profile, solving the problem of “Who-Spoke” in advance. Then, a neural network-based model takes speech features and all speaker profiles to predict their corresponding framewise voice activities, aiming to address the “When” problem. This two-stage process has demonstrated excellent performance in popular benchmarks such as DIHARD-III [21] and VoxSRC21-23 [22, 23, 24].
@@
In this article, we propose a novel MIMO-TSVAD framework for audio-visual speaker diarization in modality-missing scenarios. Let $\mathbf{X}_{a}$ and $\mathbf{X}_{v}$ denote the audio and video features, respectively. $\mathbf{E}_{spk}$ represents the target-speaker embeddings in classical TS-VAD systems. We additionally define target-lip embeddings to indicate speaker identities of corresponding lip tracks, denoted as $\mathbf{E}_{lip}$. Depending on the accessibility of different data during inference, there are four cases , namely $\mathbf{X}_{a}$ vs. $\mathbf{E}_{spk}$, $\mathbf{X}_{v}$ vs. $\mathbf{E}_{lip}$, $\mathbf{X}_{a}+\mathbf{X}_{v}$ vs. $\mathbf{E}_{lip}$, $\mathbf{X}_{a}+\mathbf{X}_{v}$ vs. $\mathbf{E}_{spk}+\mathbf{E}_{lip}$. Our proposed framework can flexibly process arbitrary cases with state-of-the-art performance and strong robustness.
Fig. 1 illustrates the progress of our TS-VAD research. The classical TS-VAD systems [17, 18] can be abstracted into the encoder-only methods, shown in Fig. 1a. $\mathbf{X}_{a}\in\mathbb{R}^{\mathrm{T\times F}}$ is extracted from input audio by traditional methods (e.g., MFCC, Fbank) or deep neural networks. $\mathbf{E}_{spk}\in\mathbb{R}^{\mathrm{N\times S}}$ represents the given speaker embeddings. Each speaker embedding $\in\mathbb{R}^{\mathrm{S}}$ has to be repeated $T$ times and concatenated with $\mathbf{X}_{a}$, producing a 3-dimensional tensor with the shape of $T\times N\times(F+S)$. Then, encoders (e.g., Bi-LSTM, Transformers) process the input tensor along the time axis ($T$) and speaker axis ($N$). As shown in Fig. 1b, we introduce the sequence-to-sequence architecture [19] to factorize the input onto the encoder and decoder separately. This way, the consumption of input memory becomes proportional to $T\times F+N\times S$, enabling the model to process longer audio and more speakers. Moreover, in the previous encoder-only architecture, the output feature length ($T^{\prime}$) must be equal to the input feature length ($T$). In the sequence-to-sequence architecture, adjusting the output dimension ($T^{\prime}$) of the last linear layer can be easily implemented. As the model accepts a fixed-length speech chunk as input when $T$ is determined, using a larger output length ($T^{\prime}$) to predict the fixed-length voice activities can achieve a higher temporal resolution. In this work, the MIMO-TSVAD framework is designed based on the sequence-to-sequence architecture, demonstrated in Fig. 1c. The multi-modal encoder and multi-task decoder leverage cross-modal and inter-speaker relationships, which can utilize various types of input data to obtain different predictions in a Multi-Input and Multi-Output (MIMO) manner.
@@ II-B End-to-End Neural Diarization
The EEND framework [8, 9] formulates the diarization problem as a multi-label classification task, relying on the permutation-invariant training to predict all speakers’ voice activities simultaneously. The initial EEND models have a fixed number of output speakers limited by the network architecture. Although the Encoder-Decoder based Attractor (EDA) [10, 11] enables EEND-based methods to process audio with a variable number of speakers, the maximum number of speakers is empirically bounded by the training data. To make the number of output speakers flexible and unlimited, EEND-vector clustering (EEND-VC) [12, 13, 14] integrates end-to-end and clustering approaches, which deploys an EEND model for shortly divided audio blocks and then matches the inter-block speaker labels by clustering on speaker embeddings. In addition, EEND-GLA [15, 16] calculates local attractors from each short block and finds the speaker correspondence based on similarities between inter-block attractors. As its training only requires relative speaker labels within the recording, EEND-GLA is practical for adapting models on in-the-wild datasets without globally unique speaker labels.
@@ II-C Target-Speaker Voice Activity Detection
The background of TS-VAD can be traced back to the personal VAD [61], which utilizes a given acoustic footprint as the speaker profile to retrieve his or her personalized voice activities. However, the personal VAD only accepts a single speaker at once and ignores the inter-speaker modeling in a conversation. Thus, TS-VAD is designed to process multiple speaker profiles and predict their voice activities simultaneously.
The initial TS-VAD [17] takes speech features (e.g., MFCC) and i-vectors of each speaker as the input, where the output number of speakers is fixed. He et al. [62] modify the model to cope with a flexible number of speakers by determining the maximum number of speakers and outputting null speech activities for zero-padded speaker profiles. Later, LSTM [63] and Transformer [20] modules are implemented along the speaker dimension of models to handle a variable number of speakers. On the other hand, the i-vectors used in TS-VAD are relatively domain-dependent, restricting the system performance on multi-scenario datasets [64]. This finding paves the way for exploring more discriminative speaker embeddings like x-vectors as an alternative. Wang et al. [18] first replace the front-end of TS-VAD with a pre-trained module tailored for extracting frame-level x-vectors. This modification shows superior robustness and generalization than a simple swap of i-vectors for x-vectors in early attempt [17].
@@
Figure 2: The MIMO-TSVAD framework. (a) Extractor: Front-end modules extract audio features $\mathbf{X}_{a}$ and video features $\mathbf{X}_{v}$ from the input data. (b): Encoder: Conformer-based modules leverage multi-modal information to generate the updated audio-visual features $\mathbf{\hat{X}}_{a}$ and $\mathbf{\hat{X}}_{v}$. (c) Decoder: Multi-task decoding modules predict voice activities based on different speaker profiles, including target-speaker embeddings $\mathbf{E}_{spk}$, target-lip embeddings $\mathbf{E}_{lip}$, or both. Green, blue, and yellow depict sub-components for processing audio, visual, and audio-visual modalities. For clarity, layer normalization and residual connection of each self-attention, cross-attention, and feedforward layer are omitted in the plot.
@@ II-D Audio-Visual Speaker Diarization
As facial attributes and lip movements have been proven to be highly related to speech [70], most early audio-visual speaker diarization methods leverage multi-modal cues by modeling the correspondence between speech signals and talking faces [25, 26]. Also, sound source localization using microphone arrays can establish another cross-modal relationship by mapping the speech direction onto the captured image plane [27, 28]. However, it is sometimes hard to perfectly enroll talking faces or lip movements with related speech segments. Off-screen speakers whose faces are not captured can lead to face detection failure. This “enroll first, diarize later” paradigm may fail in modality-missing scenarios.
Recently, Xu et al. [29] create the AVA Audio-Visual Diarization (AVA-AVD) database containing diverse movie clips and a multi-stage audio-visual speaker diarization system. The proposed method first utilizes VAD module to find speaking segments in the audio, then applies the ASD [34] module to locate a face for each speaking segment. The paired audio-visual inputs are jointly scored to cluster the speaking segments into different identities. Subsequently, Wuerkaixi et al. [30] introduce the lip movement to learn a Dynamic Vision-guided Speaker Embedding (DyViSE) instead of the still facial feature, achieving new state-of-the-art performance on these real-world videos.
In addition, the MISP 2022 [37] database is presented for Chinese Home-TV scenarios. It has over 100 hours of audio-visual signals synchronously captured by different devices, nearly 3.5 times larger than the AVA-AVD [29]. The MISP 2022 Challenge has been successfully held based on this database. Meanwhile, an End-to-End Audio-Visual Speaker Diarization (AVSD) [31] is presented as a competitive baseline method that predicts speech probabilities using audio features and each speaker’s lip video. In the AVSD system, available lip videos decide the number and order of output speakers. This way is highly effective when the camera captures all speakers ideally. In contrast, it cannot deal with the out-of-screen speakers.
@@ III-A Architecture
Fig. 2 demonstrates our proposed MIMO-TSVAD framework, consisting of three parts: the extractor, the encoder, and the decoder.
@@ III-A2 Video Extractor
[…]
@@ III-A4 Decoder
The input of each decoder branch is initialized by zero embeddings $\in\mathbb{R}^{\mathrm{N\times D}}$ and updated step by step in the subsequent blocks, where $N$ denotes the number of candidate speakers and $D$ represents the decoder dimension. The self-attention layer is adopted to exchange inter-speaker relationships. For the audio-based and video-based branches, $\mathbf{E}_{spk}$ and $\mathbf{E}_{lip}$ go through multi-layer perception (MLP) modules to align the feature dimension with the decoder dimension $D$ and concatenate with the original queries in respective branch. The adopted MLP module consists of two linear layers with in-between layer normalization and ReLU activation. Meanwhile, we concatenate keys from $\mathbf{\hat{X}}_{a}$ and $\mathbf{\hat{X}}_{v}$ with corresponding positional embeddings $\mathbf{P_{a}}$ and $\mathbf{P_{v}}$. This way, all key-query calculations in the cross-attention layer can incorporate positional and speaker-related information as the auxiliary feature. Additionally, let $\mathbf{\hat{F}}^{i}_{a}\in\mathbb{R}^{\mathrm{N\times D}}$ and $\mathbf{\hat{F}}^{i}_{v}\in\mathbb{R}^{\mathrm{N\times D}}$ denote the intermediate features of audio-based and video-based branches in the $i$-th decoder block, respectively. Then, they are concatenated to produce the audio-visual features $\mathbf{\hat{X}}_{av}\in\mathbb{R}^{\mathrm{N\times 2\times D}}$. For the mixed branch, it takes the second (modality) dimension of $\mathbf{\hat{X}}_{av}$ as the sequence axis to perform the cross-attention operation. This way, two modalities are dynamically fused into the mixed branch. Finally, a linear projection layer with sigmoid activation transforms the decoder embeddings into posterior probabilities of the estimated voice activities. The output dimension $T^{\prime}$ of the linear projection controls the temporal resolution of system prediction. For example, if the chunk length of input audio-visual data is fixed at $L_{chunk}$ seconds. The prediction for this duration will be evenly divided into $T^{\prime}$ frames. Each frame-level output indicates whether the target speaker is speaking during the corresponding time interval. In this case, the temporal resolution can be calculated as $L_{chunk}/T^{\prime}$, representing the unit duration of each frame-level prediction. By adjusting the linear projection layer, $T^{\prime}$ can be easily set as several multiples of $T_{1}$ or $T_{2}$. Given $N$ speakers, the predictions from the audio-based, video-based, and mixed output are represented as $\mathbf{\hat{Y}}_{aud}$, $\mathbf{\hat{Y}}_{vid}$, and $\mathbf{\hat{Y}}_{mix}\in\mathbb{R}^{\mathrm{N\times T^{\prime}}}$, respectively.